### v1

@@ Abstract
Targeted stimulation of just 10–20% of the most central neurons in the source population significantly enhances spiking propagation fidelity compared to random selection ($p<0.001$). This approach yields a $\sim$2.7-fold increase in signal transfer efficiency at critical inter-module connection densities ($p_{\mathrm{inter}}=0.04-0.07$). These findings establish a theoretical foundation for precision neuromodulation in biological neural systems and neurotechnology applications.
@@ 3 Results
Figure 2: Propagation efficiency in non-boosted networks. Performance surfaces show significant divergence between top-centrality (colored solid) and proxy stimulation strategies (colored shaded) and random baseline (pink solid).
@@
    Medium connectivity regime ($p_{\mathrm{inter}}=0.07$): Stimulation of top eigenvector centrality nodes achieved moderate propagation ($0.64\pm 0.24\leavevmode\nobreak\ \text{Hz}$), consistent with their role in global network integration. Strikingly, proxy stimulation targeting betweenness-central nodes near hub regions showed severely attenuated efficacy ($0.15\pm 0.14\leavevmode\nobreak\ \text{Hz}$), representing a 76% reduction versus optimal targets. This performance approached the random baseline ($0.01\pm 0.01\leavevmode\nobreak\ \text{Hz}$), demonstrating that anatomical adjacency to hubs is insufficient for effective control.
@@
    High connectivity regime ($p_{\mathrm{inter}}=0.10$): Degree-centrality targeting maximized propagation ($3.33\pm 0.98\leavevmode\nobreak\ \text{Hz}$), leveraging local integration capacity. Proxy stimulation using degree-central nodes near hubs remained markedly inefficient ($0.57\pm 0.51\leavevmode\nobreak\ \text{Hz}$, 82% reduction). The persistent strategy gap across densities suggests that conventional hub-centric approaches neglect critical control mechanisms beyond local connectivity.
Topological Enhancement Reshapes Control Landscapes Centrality boosting amplified global propagation efficiency by 2.7$\times$ ($p<10^{-10}$) and fundamentally altered network dynamics:
@@
    Threshold modulation: Boosted networks exhibited significant propagation at $p_{\mathrm{inter}}=0.07$below the non-boosted percolation threshold. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
@@
    • At $p_{\mathrm{inter}}=0.07$, closeness-centrality targeting dominated ($2.06\pm 0.44\leavevmode\nobreak\ \text{Hz}$), while proxy stimulation remained ineffective ($0.41\pm 0.38\leavevmode\nobreak\ \text{Hz}$, 80% reduction)
    • At $p_{\mathrm{inter}}=0.10$, closeness-centrality again prevailed ($7.69\pm 0.62\leavevmode\nobreak\ \text{Hz}$), but proxy stimulation showed relative improvement ($2.59\pm 1.48\leavevmode\nobreak\ \text{Hz}$, 66% reduction)
Figure 3: Strategic convergence in boosted networks. (a) Non-boosted networks exhibit distinct performance surfaces for top-centrality (solid color), proxy stimulation (shaded), and random (black) strategies. (b) Following centrality boosting, performance surfaces converge, indicating reduced efficacy differences between strategies. Shading denotes standard deviation across network ensembles.
@@ 4 Discussion
Our results demonstrate that topological centrality provides an effective criterion for driver node selection in modular neural networks. The superior performance of top-centrality nodes stems from their dual advantage: (1) high integration within the source population enables efficient input amplification, and (2) preferential connection to between-module bridges facilitates inter-population propagation.
The limited efficacy of proxy stimulation suggests that proximity to target hubs alone is insufficient for efficient control. This aligns with recent findings that effective control requires both local integration and global bridging capabilities [5]. Our spectral analysis further indicates that centrality-based selection preserves temporal fidelity of transmitted signals.
Notably, the performance gap between selection strategies was most pronounced at critical $p_{\mathrm{inter}}$ values (0.04-0.07), corresponding to the theoretical percolation threshold for modular networks. This suggests that topological control strategies are particularly valuable near phase transitions in network dynamics.
@@ 5 Conclusion
We establish that targeted stimulation of top-centrality neurons significantly outperforms both proximity-based and random selection strategies for activity propagation between neural populations. Our key findings demonstrate:
  • Critical operating regime: Centrality-based selection achieves maximum advantage ($\sim$2.7$\times$ efficiency gain) near network percolation thresholds ($p_{\mathrm{inter}}$ = 0.04-0.07), where biological neural systems typically operate
  • Metric-specific performance: Degree and closeness centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement
  • Mechanistic distinction: Spectral analysis revealed that centrality-based stimulation preserves temporal fidelity of transmitted signals, while proxy strategies amplify broadband noise
These results establish that effective neural control requires global topological influence rather than mere proximity to target hubs. The identified performance peak at critical connection densities provides a principled framework for precision neuromodulation in neurological disorders characterized by synchronization deficits. Future work should extend this approach to dynamic networks with adaptive topology and validate findings in in vivo recordings.

### v2

@@ Abstract
Targeted stimulation of just 10-20% of the most central neurons in the source population significantly enhances spiking propagation fidelity compared to random selection. This approach yields a $\sim$64-fold increase in signal transfer efficiency at critical inter-module connection densities. These findings establish a theoretical foundation for precision neuromodulation in biological neural systems and neurotechnology applications.
@@ 3 Results
Figure 2: Propagation efficiency in non-boosted networks. This figure depicts the distribution of average activity frequencies in the second network cluster following targeted stimulation of high-centrality nodes within the first cluster, specifically focusing on the top $20\%$ percentile for each centrality metric. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three data planes representing distinct stimulation conditions: the average activity response in the second cluster when stimulating the top-centrality nodes ($V_{t}$) within the first cluster; the response when stimulating proxy nodes ($V_{t}$) in the first cluster; and a random baseline response from stimulating randomly selected node subsets ($10\%,15\%,20\%$) in the first cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated.
@@
    Medium connectivity regime ($p_{\mathrm{inter}}=0.07$): Stimulation of top eigenvector centrality nodes achieved moderate propagation ($0.64\pm 0.24\leavevmode\nobreak\ \text{Hz}$), consistent with their role in global network integration. Interestingly, proxy stimulation targeting the central nodes in the kinship of the betweenness near the central regions showed severely attenuated efficacy ($0.15\pm 0.14\leavevmode\nobreak\ \text{Hz}$), representing a 4.26-fold reduction versus optimal targets. This performance approached the random baseline ($0.01\pm 0.01\leavevmode\nobreak\ \text{Hz}$), demonstrating that anatomical adjacency to hubs is insufficient for effective control.
@@
    High connectivity regime ($p_{\mathrm{inter}}=0.10$): Degree-centrality targeting maximized propagation ($3.33\pm 0.98\leavevmode\nobreak\ \text{Hz}$), leveraging local integration capacity. Proxy stimulation using degree central nodes near the hubs remained markedly inefficient ($0.57\pm 0.51\leavevmode\nobreak\ \text{Hz}$, a 5.84-fold reduction). The persistent strategy gap across densities suggests that conventional hub-centric approaches neglect critical control mechanisms beyond local connectivity.
Topological Enhancement Reshapes Control Landscapes Centrality boosting amplified global propagation efficiency and altered network dynamics:
@@
    Threshold modulation: Boosted networks exhibited significant propagation amplification, compared to non-boosted case 2.3$\times$ increase in propagation efficiency compared to non-boosted networks at $p_{\mathrm{inter}}=0.10$, with the mean firing rate in $V_{o}$ rising from $3.33\pm 0.98$ Hz to $7.69\pm 0.62$ Hz under optimal driver selection. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
@@
    • At $p_{\mathrm{inter}}=0.07$, closeness-centrality targeting dominated ($2.06\pm 0.44\leavevmode\nobreak\ \text{Hz}$), while proxy stimulation remained ineffective ($0.41\pm 0.38\leavevmode\nobreak\ \text{Hz}$, 5.02-fold reduction)
    • At $p_{\mathrm{inter}}=0.10$, closeness-centrality again prevailed ($7.69\pm 0.62\leavevmode\nobreak\ \text{Hz}$), but proxy stimulation showed relative improvement ($2.59\pm 1.48\leavevmode\nobreak\ \text{Hz}$, 2.96-fold reduction)
Figure 3: Propagation efficiency in boosted networks. This figure depicts the distribution of average activity frequencies in the second network cluster following targeted stimulation of high-centrality nodes within the first cluster, specifically focusing on the top $20\%$ percentile for each centrality metric. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three data planes representing distinct stimulation conditions: the average activity response in the second cluster when stimulating the top-centrality nodes ($V_{t}$) within the first cluster; the response when stimulating proxy nodes ($V_{t}$) in the first cluster; and a random baseline response from stimulating randomly selected node subsets ($10\%,15\%,20\%$) in the first cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated.
@@ 4 Discussion
Our results demonstrate that topological centrality provides an effective criterion for driver node selection in modular neural networks. The superior performance of top-centrality nodes stems from their dual advantage: (1) high integration within source populations enables efficient input amplification, and (2) preferential connection to inter-module bridges facilitates propagation.
The limited efficacy of proxy stimulation reveals that spatial proximity to target hubs is insufficient for effective control, aligning with recent findings that optimal control requires both local integration and global bridging capabilities [5]. Spectral analysis confirms that centrality-based selection preserves temporal fidelity of transmitted signals.
Notably, the performance gap between selection strategies was most pronounced at critical $p_{\mathrm{inter}}$ values (0.07), corresponding to theoretical percolation thresholds for modular networks. This indicates topological control strategies are particularly valuable near network phase transitions—precisely where biological neural systems typically operate.
@@ 5 Conclusion
We establish that targeted stimulation of top-centrality neurons significantly outperforms both proximity-based and random selection strategies for activity propagation between neural populations. Key findings demonstrate:
  • Critical operating regime: Centrality-based selection achieves maximum advantage (64×64$\times$64 × efficiency gain over random baseline, $p_{\mathrm{value}}<10^{-9}$) near network percolation thresholds
  • Metric-specific performance: Degree and closeness centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement (>5×>5$\times$> 5 × advantage over proxy strategies)
  • Mechanistic distinction: Centrality-based stimulation preserves temporal signal fidelity while proxy strategies amplify broadband noise
These results establish that effective neural control requires global topological influence rather than mere proximity to target hubs. The identified performance peak at critical connection densities provides a principled framework for precision neuromodulation in neurological disorders characterized by synchronization deficits. Future work should extend this approach to dynamic adaptive networks and validate findings in in vivo recordings.